### v2

@@ 4 Methodology
To effectively fine-tune xFinder, a high-quality dataset is essential. To our knowledge, there is currently no dedicated dataset for Key Answer Extraction. In this section, we will detail the implementation process of xFinder, including the construction of the KAF dataset and the training of xFinder, as illustrated in Figure 3. The KAF dataset encompasses a variety of evaluation tasks, including questions, optional answer ranges, LLM responses to the questions, and the extracted key answers. This dataset is divided into three segments: a training set, a test set, and a generalization set, which are used for fine-tuning, testing, and assessing xFinder’s performance, respectively. The training set comprises 26,900 samples, the test set includes 4,961 samples, and the generalization set contains 4,482 samples. For more detailed information, please refer to the Appendix B.1. Section 4.1 details the process of LLM response collection, Section 4.2 describes our method and workflow for annotating key answers in the LLM-generated content, and Section 4.3 elaborates on our detailed approach to training xFinder.
Figure 3: Schematic of the research framework. The first three stages correspond to Sections 4.1,  4.2, and  4.3 respectively; the final stage illustrates the replacement of RegEx with xFinder in the current evaluation pipeline. The experiments in Section 5.2 demonstrate the efficacy of our approach within this pipeline. Note: The percentages 80.3%, 77.0%, and 2.2% in the center of the figure around ’unreliable evaluation’ indicate results from the Llama3-8B-Instruct on the GSM8K benchmark using RegEx evaluation via the LM Eval Harness, OpenCompass, and UltraEval frameworks, respectively, while our method achieves a reliable result of 80.2%.
@@ 4.1 LLM Response Generation
As shown in the first part of Figure 3, we initially generate question-response pairs using different LLMs across multiple datasets. Currently, the tasks for evaluating LLM performance are quite diverse. We investigated the leading LLM benchmarks and evaluation reports to cover a wide range of popular evaluation tasks. Ultimately, we selected 18 typical evaluation task datasets, with detailed information provided in Appendix A.1. We categorized these tasks into four types, each with similar option formats, which serve as prompts to induce the LLMs to generate varied responses (as shown in the left boxes in Figure 2): (1) Alphabet option tasks. By referring to the design of mainstream evaluation rankings, these tasks are formatted as alphabet options, where each task aims to select the single correct option from multiple possibilities (i.e., A, B, C, D, etc.). (2) Short text option tasks. The requirement for the model to correctly map letters to option texts can affect the assessment of the model’s true capabilities [36]. Therefore, we designed tasks with key answer types as short texts in the form of alphabet options, where the goal is to select the single correct short text from multiple short texts. (3) Categorical label tasks. For traditional classification tasks (e.g., sentiment classification, topic classification, entity classification), we retained their original format, using the labels themselves as options. (4) Math option tasks. Given the output types of mathematical tasks, which could be integers, decimals, or even LaTeX code, we categorized them separately.
To further enhance the heterogeneity and richness of the dataset, we have primarily designed eight types of prompt configurations to elicit diverse responses from the LLMs. These configurations include 0-shot(-restrict444-restrict indicates that the prompt ends with the sentence ”End your final answer with ’The answer is ¡answer¿.’” to standardize the LLM’s response format.), 0-shot-cot(-restrict), 5-shot(-restrict), and 5-shot-cot(-restrict). Specific examples of these prompts are provided in Appendix F.1.
For generating responses, it is crucial to capture a wide range of output content, given that LLMs of different series and sizes exhibit varied reasoning processes and may provide different answers to the same question. Consequently, we utilized 10 distinct LLMs for this purpose. Specific details on these models can be found in Appendix A.2.
It is important to note that the KAF dataset’s training and test sets consist of Question-Response Pairs generated by 10 LLMs across 8 evaluation tasks. To ensure the validity and fairness of the generalization evaluation, the Question-Response Pairs used in the generalization set are derived from responses by 8 LLMs on 11 entirely new evaluation tasks.
@@ 4.2 Auto Labelling and Human Recheck
The foundation of constructing the KAF dataset is the extraction of the key answer $\boldsymbol{k}$ from the LLM response $\boldsymbol{y}$, which is then used as the label. We employed distinct annotation strategies for the training, test, and generalization sets. For the training set, to improve the efficiency of the annotation process, a semi-automated procedure was adopted. This involved generating two sets of annotations from GPT-4 using varied prompts F.2. Utilizing the Self-Consistency strategy [37], items with discrepancies between the two annotation rounds, along with all mathematics questions, were consolidated for manual re-annotation using Label Studio [38]. To ensure the validity and reliability of the test and generalization sets, all labels needed to be accurate, necessitating two rounds of manual annotation for these datasets. Detailed information about the annotation methods can be found in Appendix B.2.
@@ 4.3 Training xFinder
According to the definition of key answer extraction outlined in Section 3, our goal is to fine-tune xFinder to precisely locate $\boldsymbol{s}$ within $\boldsymbol{y}$ and accurately extract $\boldsymbol{k}$ from $\boldsymbol{s}$. The training input for xFinder includes the question $\boldsymbol{q}$, the LLM response $\boldsymbol{y}$, the key answer type, and the answer range, while the output is the key answer $\boldsymbol{k}$.
During the training process, we primarily utilized the XTuner [39] tool developed by the InternLM team for fine-tuning. To fully exploit the instruction-following potential of xFinder, we meticulously crafted prompts as demonstrated in Appendix F.3. Additionally, to enhance the generalization capability of xFinder, we implemented data augmentation on the KAF training set, which included:
@@
  • Enriching the variety of prompt forms in Prompt wrapped key answer sentences. From the LLM responses in the KAF training set that contained wrapped key answer sentences, we extracted 10% and substituted the $\boldsymbol{f}$ with other elements from $\mathcal{F}$, such as replacing "The final answer is A" with "Based on the context of the question, A is the most likely answer."
@@ 5 Experiments
In this section, we initially fine-tuned 19 LLMs with parameter sizes ranging from 0.5B to 8B using the KAF training set. We compared their accuracy in extracting key answers on both the KAF test and generalization sets against the mainstream evaluation frameworks using RegEx methods. Subsequently, we evaluated the performance of 10 LLMs on 14 evaluation task datasets using xFinder-qwen1505, xFinder-llama38it, and the current mainstream evaluation frameworks, analyzing differences in evaluation results.
@@
We employed the QLoRA method from the XTuner framework [39] [40] for fine-tuning the foundation models, with specific tuning parameters detailed in Appendix C. To assess the efficacy of xFinder, we conducted tests using the KAF test set. We set the current mainstream frameworks—OpenCompass, LM Eval Harness, and UltraEval—with their RegEX extraction modules as baselines. The comparison metric is extraction accuracy, defined as the proportion of correctly extracted key answers relative to the dataset size. Partial results are presented in Table 1; for complete experimental results, refer to Table 18 in Appendix D.1.
alphabet option short text categorical label math Overall
@@
Table 1: Comparison of Extraction Accuracy on KAF Test Set: Mainstream Frameworks vs. xFinder Models. Specifically, xFinder-qwen1505, xFinder-qwen1518, xFinder-gemma7, xFinder-chatglm36base, and xFinder-llama38 are models fine-tuned using the KAF training set on Qwen1.5-0.5B, Qwen1.5-1.8B, Gemma-7B, ChatGLM3-6B-base, and Llama-8B, respectively.
It is evident that the xFinder models, fine-tuned using the KAF training set, are highly effective. Furthermore, despite a 16-fold difference in the parameter sizes of the models, the variation in extraction accuracy does not exceed 0.5%. This implies the high quality of the KAF dataset we constructed.
@@
To further explore xFinder’s extraction capabilities on not-seen tasks, we conducted tests using the KAF generalization set. Specific results are shown in Table 2. Only two xFinder models are displayed here for brevity, but the complete experimental results can be found in Table 19 of Appendix D.1. The results indicate that xFinder maintains a high level of extraction accuracy on the generalization set. In contrast, mainstream regular expression extractors perform poorly [^5] Here, the extraction accuracy of OpenCompass, LM Eval Harness, and UltraEval has significantly decreased compared to Table 1. This is because the LLM responses in the KAF test set are generated from restricted prompts, making it easier for regular expressions to extract the key answer; however, the responses in the KAF generalization set are generated from non-restricted prompts., especially in extracting key answers for mathematical problems. Even GPT-4 struggles to extract key answers from complex and varied LLM responses accurately.
alphabet option short text categorical label math Overall $\Delta_{acc}$ $\Delta_{acc}/N$
@@
Table 2: Comparison of Extraction Accuracy on KAF Generalization Set: Mainstream Frameworks, GPT-4, and xFinder Models. Specifically, xFinder-qwen1505 and xFinder-llama38it represent models fine-tuned with the KAF dataset on Qwen1.5-0.5B and Llama3-8b-instruct, respectively.
@@ 5.2 Evaluation in Real-World Scenarios
The experiments conclusively demonstrate that xFinder plays a critical role within evaluation frameworks. Next, we will use xFinder for automated evaluations in real-world evaluation tasks instead of only inspecting the extraction accuracy. To balance evaluation efficiency and quality, we calculated two metrics for all xFinder models on the KAF Generalization Set: $\Delta_{acc}$ and $\Delta_{acc}/N$ (where $N$ is the parameter size of the LLM, measured in billions). Based on these metrics, xFinder-qwen1505 (with the highest $\Delta_{acc}/N$) and xFinder-llama38it (with the highest $\Delta_{acc}$) were selected. Here, $\Delta_{acc}$ represents the difference in extraction accuracy on the KAF Generalization Set between the current xFinder and the lowest extraction accuracy among the 19 xFinder models. For specific results, refer to Table 19 in Appendix D.1.
We evaluated the performance of 10 LLMs across 14 mainstream evaluation tasks using xFinder-qwen1505 and xFinder-llama38it, along with regular extraction methods based on OpenCompass, LM Eval Harness, and UltraEval. The comprehensive scoring results are provided in Appendix D.2.
@@
RQ1: Are existing evaluation frameworks reliable? The graph reveals significant discrepancies in the rankings of LLMs obtained using different evaluation frameworks. For instance, in the CommonsenseQA (alphabet) task, InternLM2-1.8B-Chat ranks first in UltraEval but only fifth in OpenCompass. Such disparities in rankings indicate the unreliability of RegEx methods within current mainstream evaluation frameworks, potentially affecting model evaluation and ranking.
@@
6 Limitations
Although xFinder has achieved encouraging results, we must acknowledge some limitations. Our smallest xFinder model, with only 500 million parameters, performs well as a key answer extraction component in evaluation tasks. Nevertheless, the deployment cost may still be high for some researchers. In the future, we promise to continuously update xFinder by exploring model compression techniques to support smaller versions, potentially enabling CPU deployment and thereby further enhancing xFinder’s flexibility.

### v3

@@ 4 Methodology
To fine-tune xFinder, a high-quality dataset is essential. To our knowledge, there is no dedicated dataset for key answer extraction and matching. In this section, we will detail the implementation of xFinder, including the KAF dataset construction and model training, as illustrated in Figure 3. The KAF dataset encompasses a variety of evaluation tasks, including questions, optional answer ranges, LLM responses to the questions, and the extracted key answers. The dataset is divided into three segments: training, test, and generalization sets, used for fine-tuning, testing, and performance assessment, respectively. The training set has 26,900 samples, the test set 4,961 samples, and the generalization set 4,482 samples. For more detailed information, refer to Appendices B.1 and B.3. Section 4.1 details LLM responses collection, Section 4.2 describes the annotation of key answers in LLM-generated content, and Section 4.3 elaborates on our detailed approach to training xFinder.
Figure 3: Schematic of the research framework. The first three stages correspond to Sections 4.1,  4.2, and  4.3, while the final stage illustrates the replacement of RegEx with xFinder in the evaluation pipeline. The experiments in Section 5.3 demonstrate the efficacy of our approach within this pipeline. Note: The percentages 80.3%, 77.0%, and 2.2% in the center of the figure around “unreliable evaluation” indicate results from the Llama3-8B-Instruct on the GSM8K benchmark using RegEx evaluation via the LM Eval Harness, OpenCompass, and UltraEval frameworks, respectively, while our method achieves a reliable result of 80.2%.
@@ 4.1 LLM Response Generation
As shown in the first part of Figure 3, we initially generate question-response pairs using different LLMs across multiple datasets. Currently, the tasks for evaluating LLM performance are highly diverse.We investigated the leading LLM benchmarks and evaluation reports to cover popular evaluation tasks. Ultimately, we selected 19 typical evaluation task datasets, with detailed information provided in Appendix A.1. We categorized these tasks into four types, each with similar option formats serving as prompts to induce LLMs to generate varied responses (as shown in the left boxes in Figure 2): (1) Alphabet option tasks. By referring to the design of mainstream evaluation rankings, these tasks are formatted as alphabet options, where each task aims to select the single correct option from multiple possibilities (i.e., A, B, C, D, etc.). (2) Short text option tasks. The requirement for the model to map letters to option texts may impact the evaluation of the model’s true capabilities (Wang et al., 2024b). Therefore, we designed tasks with key answer types as short texts, where the goal is to select the correct option from multiple short texts. (3) Categorical label tasks. For traditional classification tasks (e.g., sentiment classification, topic classification, entity classification), we retained their original format, using the labels themselves as options. (4) Math option tasks. Given the variety of output types in math tasks, such as integers, decimals, and LaTeX code, we categorized them separately.
To enhance the diversity of the dataset, we designed eight types of prompt configurations to elicit varied responses from the LLMs. These configurations include 0-shot(-restrict444-restrict indicates that the prompt ends with the sentence “End your final answer with ‘The answer is ¡answer¿.’ ” to standardize the LLM’s response format.), 0-shot-cot(-restrict), 5-shot(-restrict), and 5-shot-cot(-restrict). These prompt templates are provided in Appendix F.1.
For generating responses, it is crucial to capture a wide range of output content, given that LLMs of different series and sizes exhibit varied reasoning processes and may provide different answers to the same question. Consequently, we utilized a total of more than 10 distinct LLMs, and specific details about these models can be found in Appendix A.2. The KAF dataset’s training and test sets include question-response pairs generated by 10 LLMs across 8 evaluation tasks. To ensure the validity and fairness of the generalization evaluation, the question-response pairs used in the generalization set are derived from responses by 8 LLMs on 11 entirely new evaluation tasks. Further details on the KAF dataset can be found in Appendix B.1.
@@ 4.2 Auto Labelling and Human Recheck
The foundation of the KAF dataset is extracting the key answer $\bm{k}$ from the LLM response $\bm{y}$, which serves as the label. We employed distinct annotation strategies for the training, test, and generalization sets. For the training set, to improve annotation efficiency, we adopted a semi-automated procedure. This involves generating two sets of annotations from GPT-4 with different prompts (refer to Appendix F.2 for prompt templates). Using the self-consistency strategy (Wang et al., 2023a; Liang et al., 2024b), items with discrepancies between the two annotation rounds, along with all mathematics questions, were manually re-annotated using Label Studio (Tkachenko et al., 2020). Two rounds of manual labeling were conducted to ensure the accuracy and reliability of the test and generalization sets. Detailed information about the annotation methods can be found in Appendix B.2.
@@ 4.3 Training xFinder
According to the definition of key answer extraction outlined in Section 3, our goal is to fine-tune xFinder to precisely locate $\bm{s}$ within $\bm{y}$ and accurately extract $\bm{k}$ from $\bm{s}$. The training input for xFinder includes the question $\bm{q}$, the LLM response $\bm{y}$, the key answer type, and the answer range, while the output is the key answer $\bm{k}$.
During the training process, we primarily utilized the XTuner (InternLM, 2023) tool developed by the InternLM team for fine-tuning. To fully exploit the instruction-following potential of xFinder, we meticulously crafted prompts as demonstrated in Appendix F.4. Additionally, to enhance the generalization capability of xFinder, we implemented data augmentation on the KAF training set, which included:
@@
  • Enriching the variety of prompt forms in Prompt wrapped key answer sentences. From the LLM responses in the KAF training set that contained wrapped key answer sentences, we extracted 10% and substituted the $\bm{f}$ with other elements from $\mathcal{F}$, such as replacing “The final answer is A” with “Based on the context of the question, A is the most likely answer.”
@@ 5 Experiments
In this section, we first outline the experimental setup, including metrics, baselines, and fine-tuning configurations, which apply to the Extraction Accuracy and Judgment Accuracy experiments.
  • Metrics. For the Extraction Accuracy and Judgment Accuracy experiments, we employ two primary metrics: Extraction Accuracy, measuring the proportion of correctly extracted key answers, and Judgment Accuracy, evaluating the correctness of judgments on LLM responses. Detailed definitions of these metrics are provided in Appendix D.1.
  • Baselines. The baselines for these two experiments include mainstream LLM evaluation frameworks such as OpenCompass (OpenMMLab, 2023), LM Eval Harness (Gao et al., 2021), and UltraEval (He et al., 2024), all of which rely on RegEx-based extraction methods. Additionally, we compare with judge models such as PandaLM (Wang et al., 2023b) and JudgeLM (Zhu et al., 2023), as well as GPT-4, which serves as both an extractor and a judge (the prompt used is provided in Appendix F.2 and Appendix F.3).
  • Fine-tuning Details. For the experiments, we fine-tuned 19 LLMs, with parameter sizes ranging from 0.5B to 8B, using the KAF training set. Fine-tuning was performed using the QLoRA method from the XTuner framework (InternLM, 2023; Dettmers et al., 2024), with detailed parameters provided in Appendix C.
Finally, for the Real-world Evaluation experiment, we evaluated the performance of 10 LLMs on 14 evaluation task datasets, comparing results obtained using xFinder-qwen1505, xFinder-llama38it, and the mainstream evaluation frameworks. This experiment analyzes differences in evaluation outcomes to establish xFinder’s reliability in practical applications.
@@
We employed the QLoRA method from the XTuner framework (InternLM, 2023; Dettmers et al., 2024) for fine-tuning the foundation models, with specific tuning parameters detailed in Appendix C. To evaluate the effectiveness of xFinder, we conducted tests using the KAF test set. We set the mainstream frameworks—OpenCompass, LM Eval Harness, and UltraEval—with their RegEX extraction modules as baselines. The comparison metric is extraction accuracy, defined as the proportion of correctly extracted key answers relative to the dataset size. Partial results are presented in Table 1; For complete experimental results, refer to Table 22 in Appendix D.2.
Table 1: Comparison of Extraction Accuracy on KAF Test Set: Mainstream frameworks vs. xFinder models. Specifically, xFinder-qwen1505 and xFinder-llama38it represent models fine-tuned with the KAF training set on Qwen1.5-0.5B and Llama3-8B-Instruct, respectively.
Method alphabet option short text categorical label math Overall
@@
It is evident that the xFinder models, fine-tuned using the KAF training set, are highly effective. Furthermore, despite a 16-fold difference in the parameter sizes of the models, the variation in extraction accuracy does not exceed 0.5%.
@@
To further explore xFinder’s extraction capabilities on not-seen tasks, we conducted tests using the KAF generalization set. Specific results are shown in Table 2. Only two xFinder models are displayed here for brevity, but the complete experimental results can be found in Table 23 of Appendix D.2. The results indicate that xFinder maintains a high level of extraction accuracy on the generalization set. In contrast, mainstream RegEx methods perform poorly [^5] Here, the extraction accuracy of OpenCompass, LM Eval Harness, and UltraEval has significantly decreased compared to Table 1. This is because the LLM responses in the KAF test set are generated from restricted prompts, making it easier for RegEx methods to extract the key answer; However, the responses in the KAF generalization set are generated from non-restricted prompts., especially in extracting key answers for mathematical problems. Even GPT-4 struggles to extract key answers from complex and varied LLM responses accurately.
Table 2: Comparison of Extraction Accuracy on KAF Generalization Set: Mainstream frameworks, GPT-4, and xFinder models. Specifically, xFinder-qwen1505 and xFinder-llama38it represent models fine-tuned with the KAF training set on Qwen1.5-0.5B and Llama3-8b-instruct, respectively.
Method alphabet option short text categorical label math Overall $\Delta_{acc}$ $\Delta_{acc}/N$
@@
5.2 Judgment Accuracy: xFinder vs. Judge Models
As shown in Table 3, we present the evaluation results of mainstream frameworks, judge models, GPT-4, and xFinder on the generalization set. The metric is Judgment Accuracy, which measures the correctness of judgments on LLM responses. We observed that the judgment accuracy of PandaLM was only 51.9%, while JudgeLM-33B achieved merely 78.13%. In addition, the highest accuracy of GPT-4’s judgments is only 88.42%, which is significantly lower than that of xFinder.
Table 3: Comparison of Judgment Accuracy on KAF Generalization Set: Mainstream frameworks, judge models, GPT-4, and xFinder, using human-annotated results as ground truth. All judge models use prompts with reference answers.
Method alphabet option short text categorical label math Overall
OpenCompass 0.8742 / / 0.9125 0.8870
LM Eval Harness 0.8117 0.9148 0.9750 0.5813 0.8592
UltraEval 0.7836 / / 0.5328 0.7000
PandaLM-7B 0.4953 0.5832 0.5312 0.4391 0.5190
JudgeLM-7B 0.7195 0.8316 0.8056 0.5875 0.7555
JudgeLM-13B 0.6875 0.8545 0.7694 0.9266 0.7867
JudgeLM-33B 0.8133 0.8358 0.6906 0.8625 0.7813
GPT-4 as Judge 0.9016 0.8909 0.7294 0.9313 0.8420
GPT-4 as Judge (CoT) 0.9234 0.9345 0.7919 0.9609 0.8842
xFinder-qwen1505 0.9781 0.9761 0.9625 0.9969 0.9748
xFinder-llama38it 0.9750 0.9688 0.9731 0.9969 0.9761
We present the differences between answer Extraction Accuracy and final Judgment Accuracy in Table 4. From the results, we can see that the baseline method with the smallest difference, OpenCompass, still has a judgment accuracy 14.32% higher than its extraction accuracy, indicating a high unreliability in the evaluation results obtained using traditional RegEx-based key answer extraction methods. In contrast, the judgment accuracy of our xFinder-llama38it method is only 2.43% higher than its extraction accuracy, reducing judging reliability due to inaccurate extraction to within 3%.
Table 4: The difference, that is Accuracy Gap ($\downarrow$), between Judgment Accuracy ($\uparrow$) and Extraction Accuracy ($\uparrow$) on the Generalization Set for Mainstream frameworks, GPT-4, and xFinder. It reflects discrepancies in Judgment Accuracy due to extraction errors. This set uses human-annotated results as the ground truth.
Method Extraction Accuracy $\uparrow$ Judgment Accuracy $\uparrow$ Accuracy Gap $\downarrow$
OpenCompass 0.7438 0.8870 0.1432
LM Eval Harness 0.6780 0.8592 0.1812
UltraEval 0.3978 0.7000 0.3022
GPT-4 0.6957 0.8420 0.1463
xFinder-qwen1505 0.9342 0.9748 0.0406
xFinder-llama38it 0.9518 0.9761 0.0243
Efficiency Comparison
To further compare our xFinder with existing LLM as judge models, including JudgeLM, PandaLM, and the API-based GPT-4, we conducted a cost analysis experiment. We randomly sampled 200 instances from four question types within the Generalization Set. As shown in Table 5, we present the time required to evaluate these tasks on the same machine equipped with 8 NVIDIA H100 (80G) GPUs. Notably, JudgeLM-33B utilized 2 GPUs, while the other models operated on a single GPU, with all evaluations conducted in a single process. In Table 6, we present the complete evaluation tasks along with the costs incurred by the API-based GPT-4. The cost of xFinder-qwen1505 under the aforementioned setup is merely \$0.02, accounting for only 0.4% of GPT-4’s expenses. It can be observed that our xFinder achieves not only higher accuracy compared to these judge models but also lower evaluation costs and greater flexibility.
Table 5: Time costs (in seconds) for xFinder and judge models in evaluating 200 randomly selected questions per task type from KAF generalization set.
Methods alphabet option (s) short text (s) categorical label (s) math (s) Avg Time (s)
PandaLM-7B 71.05 70.09 56.45 38.88 59.12
JudgeLM-7B 228.11 227.83 240.54 330.40 256.72
JudgeLM-13B 395.57 457.49 415.46 415.31 420.96
JudgeLM-33B 522.05 527.63 517.25 571.82 534.69
xFinder-qwen1505 10.24 11.12 10.05 11.28 10.67
xFinder-llama38it 13.43 16.79 12.79 16.80 14.95
Table 6: Total costs (in USD) for GPT-4 API in evaluating 200 randomly selected questions per task type from KAF generalization set.
Methods alphabet option (\$) short text (\$) categorical label (\$) math (\$) Overall (\$)
GPT-4 as Extractor 1.34 1.2 1.13 1.39 5.06
GPT-4 as Judge 1.25 1.14 1.19 1.57 5.15
@@ 5.3 Evaluation in Real-World Scenarios
The experiments conclusively demonstrate that xFinder plays a critical role within evaluation frameworks. Next, we will use xFinder for automated evaluations in real-world evaluation tasks instead of only inspecting the extraction accuracy. To balance efficiency and quality, we calculated two metrics for all xFinder models on the KAF generalization set: $\Delta_{acc}$ and $\Delta_{acc}/N$ (where $N$ is the parameter size of the LLM, measured in billions). Based on these metrics, xFinder-qwen1505 (with the highest $\Delta_{acc}/N$) and xFinder-llama38it (with the highest $\Delta_{acc}$) were selected. Here, $\Delta_{acc}$ represents the difference in extraction accuracy on the KAF generalization set between the current xFinder and the lowest accuracy among the 19 xFinder models. For specific results, refer to Table 23 in Appendix D.2.
We evaluated the performance of 10 LLMs across 14 mainstream evaluation tasks using xFinder-qwen1505 and xFinder-llama38it, along with RegEx methods based on OpenCompass, LM Eval Harness, and UltraEval. The comprehensive scoring results are provided in Appendix D.3.
@@
RQ1: Are existing evaluation frameworks reliable? The graph reveals significant discrepancies in the rankings of LLMs obtained using different evaluation frameworks. For instance, in the CommonsenseQA (alphabet) task, InternLM2-1.8B-Chat ranks first in UltraEval but only fifth in OpenCompass. Such disparities in rankings indicate the unreliability of RegEx methods within mainstream evaluation frameworks, potentially affecting the credibility and ranking of model evaluation.